OpenStudy (anonymous):

I will medal and fan anyone who answers this: there is an isocelles triangle the base is s and the other sides are r. one of the angles is 45 the others are 67.5 Prove that s^2 = r^2 (2-sqrt2)

OpenStudy (anonymous):



OpenStudy (anonymous):

this triangle is part of an octagon and teh octagon is inside a circle

OpenStudy (anonymous):

but the triangle is the main part i gues

OpenStudy (anonymous):

Yes, I don't think we need the other geometric info to figure this out. Have you tried the law of sines?

OpenStudy (anonymous):

yes but to no avail but it might work im not sure

OpenStudy (anonymous):

\[\frac{\sin67.5^\circ}{r}=\frac{\sin45^\circ}{s}\] This identity will help later on: \[\sin^2x=\frac{1-\cos2x}{2}\]

OpenStudy (anonymous):

okay but how?

OpenStudy (anonymous):

Replace any expressions with values you know (i.e. \(\sin45^\circ\)). Then square both sides.

OpenStudy (anonymous):

^Not the only way to do it, just a suggestion.

OpenStudy (anonymous):

i tried but could not do it pls help

OpenStudy (anonymous):

@SithsAndGiggles pls help!!